ADDITION OF WATER TO AMMONIA-HYDROGEN HALIDE COMPLEXES: MOVING TOWARDS PRO-TON TRANSF ER

\underline{BRIAN E. APPLEGATE}, JEREMY M. MERRITT, AND ROGER E. MILLER, University of North Carolina at Chapel Hill, Deptartment of Chemistry, Chapel Hill, NC 27599.

Proton transfer in the ammonia-hydrogen halides serves as a prototypical acid-base reaction. Understanding the role of water in facilitating these types of reactions is important for both atmospheric chemistry and solution chemistry. Gasphase and matrix experiments^{a} show that the proton remains attached to the halogen in anhydrous complexes for HF, HCl, and HBr. However, recent theoretical calculations^{b} have shown that the proton transfer proceeds with the addition of from one to three water molecules, for HBr, HCl, and HF, respectively. To study this reaction we have initiated a set of experiments where we sequentially add water molecules to preformed ammonia-hydrogen halide complexes inside helium droplets. Observation of the vibrational frequencies along with estimates of the dipole moments help to determine if proton transfer has taken place.

aBarnes, A.; Legon, A., J. Mol. Struct., 1998, 101.

bSnyder, J.; Cazar, R.; Jamka, A.; Tao, F-M, J. Phys. Chem., 1999, 7719.